It is required to prepare 1250kg of a solution composed of 12 wt. % ethanol and 88 wt. % water. Two solutions are available, the first contains 5 wt. % ethanol, and the second contains 25 wt. % ethanol. How much of each solution are mixed to prepare the desired solution? Check your answer
1
Expert's answer
2019-09-05T03:43:47-0400

1.Ethanol balance

"A(5\/100)+B(25\/100)=M(12\/100)"

"0.05A+0.25B=0.12M"

"A=(150-0.25B)\/0.05=3000-5B"

2.Water balance

"0.95A+0.75B=0.88M=0.88*1250=1100"

"0.95A+0.75B=1100"

Sub.(1) in (2)

"0.95(300-5B)+0.75B=1100"

"2850-4.75+0.75B=1100"

"4B=1750"

"B=437.5kg"

Sub B in (1):

"A=3000-5*437.5=812.5kg"

Answer:812.5kg 5% and 437.5kg 25% ethanol